# Bye bye, local bias: the statistics of the halo field are not determined by the local mass density

### v1

@@ Abstract
Context. Bias models which relate the dark matter field to the spatial distribution of halos are widely used in current cosmological analyses. Many such models predict halos purely from the local matter density, an assumption which has not been verified in a model-agnostic setting. Bias models in perturbation theory require the inclusion of other local properties, but it is not clear whether this extends to non-perturbative approaches.
Aims. We assess the validity of the assumption that only the local dark matter density can be used to predict the number density of halos in a model-independent way and in the non-perturbative regime.
Methods. Utilising $N$-body simulations, we introduce a test wherein we study the properties of the halo counts field after spatial voxels with near-equal dark matter density have been permuted. If local-in-matter-density biasing were valid, the statistical properties of the permuted and un-permuted fields would be indistinguishable since both represent equally fair draws of the stochastic biasing model.
Results. For voxels of side length $\sim 4-60\,h^{-1}{\rm\,Mpc}$ and for halos less massive than $\sim 10^{15}\,h^{-1}{\rm\,M_{\sun}}$, we find that the permuted halo field has significantly too much power on large scales compared to the un-permuted field. We interpret this as due to these bias models removing small-scale power by not modelling correlations between neighbouring voxels. Since the permutation conserves the total variance of the halo counts field, large-scale power is substantially boosted to compensate for the missing small-scale power. This conclusion is robust to the choice of initial conditions and cosmology.
Conclusions. The assumption of local-in-matter-density halo biasing cannot, therefore, reproduce the distribution of halos across a large range of scales and halo masses, no matter how complex the model. To reproduce this distribution accurately, one must either allow the biasing to be a function of other quantities than the local matter density and/or remove the assumption that neighbouring voxels are statistically independent.
Key Words.:
Cosmology: theory, large-scale structure of Universe
@@ 1 Introduction
On the largest scales, one expects that the bias is linear, namely the overdensity of tracers is proportional to the overdensity of matter. This approximation breaks down as one moves from linear to quasilinear scales, and more general functional relationships must be used. The simplest assumption one can make is that the tracer number density depends solely on the local dark matter density - the so-called “local-in-matter-density” (LIMD) assumption (Desjacques et al., 2018) which, in a perturbative analysis, allows one to write the tracer number density as an expansion in the overdensity: the local bias expansion (Fry & Gaztanaga, 1993). More generally, one can express the bias model as local in gravitational observables, namely not just the density field but other terms related to the tidal field or higher derivative terms (Mirbabayi et al., 2015; Desjacques et al., 2018). One can enumerate such terms based on the order of perturbation theory such that, if one goes to sufficiently high order and is in the large-scale, perturbative regime, then the result converges to the truth.
@@
Relying on perturbation theory (Bernardeau et al., 2002) or effective field theory (EFT) (Carrasco et al., 2012; Senatore & Zaldarriaga, 2014; Senatore, 2015; Perko et al., 2016) inherently limits our analysis to these quasi-linear scales. Although it is possible to study higher order statistics (D’Amico et al., 2022) or perform field-level analyses in these frameworks (Schmidt et al., 2019; Elsner et al., 2020; Kostić et al., 2023; Stadler et al., 2023), to extract information on non-linear scales one may need to go beyond such local bias expansions, although some studies have shown that perturbative methods perform well when tested against $N$-body simulations (Roth & Porciani, 2011; Schmittfull et al., 2019). Instead of relying of power series, many parametric forms for the biasing relation have been proposed in the literature based on the LIMD assumption (Szalay, 1988; Matsubara, 1995, 2011; Frusciante & Sheth, 2012; Neyrinck, 2014; Ata et al., 2015). As such, many reconstruction analyses typically assume that the number density of halos is solely a function of the smoothed density field (e.g. Schmoldt et al., 1999; Erdoǧdu et al., 2004, 2006; Kitaura et al., 2009; Jasche et al., 2010; Jasche & Wandelt, 2012, 2013; Jasche et al., 2015; Lavaux & Jasche, 2016; Modi et al., 2018; Jasche & Lavaux, 2019; Lavaux et al., 2019; Ramanah et al., 2019; Kitaura et al., 2021). Alternatively, one can utilise machine learning methods to map the dark matter density field (using both local and non-local information) to the halo counts field (Charnock et al. (2020); see Dai & Seljak (2021) for a Lagrangian approach).
@@
Unlike in the perturbative regime, we do not have theoretical guarantees that any of these bias models are accurate and so this must be tested empirically. Mirbabayi et al. (2015) argue that, at any order in perturbation theory, the bias can be expressed in terms of locally measurable observables, and that this argument is closed under renormalisation. However, it could be (as is often assumed (Szalay, 1988; Matsubara, 1995, 2011; Frusciante & Sheth, 2012; Neyrinck, 2014; Ata et al., 2015)) that the only important observable is the local dark matter density field. Testing this assumption is the subject of this letter. There is an important distinction, however, between testing the assumptions of the halo biasing model (i.e. which variables are needed to predict the tracer number density) and the particular model chosen (i.e. what is the functional form of the bias or the stochasticity in the tracer-matter relation). In this letter we study the former issue in the fully non-linear regime.
@@
We present a model-independent test of the assumption of LIMD halo biasing. Our test is independent of the functional form of the biasing model, but simply tests whether halo biasing can be a stochastic function of only the local matter density. We find that LIMD halo biasing cannot reproduce the distribution of halos less massive than $\sim 10^{15}\,h^{-1}\,{\rm M_{\sun}}$ when applied to gridded dark matter fields with voxels of size less than $\sim 60\,h^{-1}{\rm\,Mpc}$. We demonstrate that this conclusion is robust across initial conditions and cosmology by utilising the quijote suite of $N$-body simulations. We conclude that no LIMD biasing model, however complex, should be used if one wishes to accurately predict the spatial distribution of halos given a matter density field.
@@
In Section 2 we explicitly define LIMD biasing, introduce our model-independent test and describe the simulations used to perform this test. Our results are presented in Section 3, where we demonstrate that our conclusions are highly robust to the one free parameter in our test, as well as cosmology and initial conditions, and investigate the adequacy of LIMD biasing as a function of scale and halo mass. Section 4 discusses these results and presents our conclusions.
@@ 2.1 Assessing the assumption of local biasing
Since the dark matter density field takes continuous values, instead of performing permutations on voxels which have exactly equal dark matter density, we first sort the voxels into $n_{\rm bin}$ bins of dark matter density, and consider two voxels equivalent if they fall within the same bin, namely if their density is approximately equal. In our fiducial analysis, we choose to use adaptive bin widths so that each bin has approximately the same number of voxels and use $n_{\rm bin}=1000$. We demonstrate that our results are highly robust to this choice in Section 3.
@@ 2.2 Simulations
For our fiducial analysis, we consider simulation number 0 from the fiducial cosmology, which matches the Planck 2018 cosmological parameters (Planck Collaboration, 2020): $\Omega_{\rm m}=0.3175$, $\Omega_{\rm b}=0.049$, $h=0.6711$, $n_{\rm s}=0.9624$, $\sigma_{8}=0.834$. In Section 3 we verify that our conclusions are robust if we chose different initial conditions (i.e. for different simulation numbers in the suite). We also study the sensitivity of our conclusions to the choice of cosmological parameters by performing our analysis with the (high-resolution) Latin-hypercube suite of quijote simulations, consisting of 2000 dark-matter-only simulations with cosmological parameters arranged on a Latin hypercube in the range: $\Omega_{\rm m}\in[0.1,0.5]$, $\Omega_{\rm b}\in[0.03,0.07]$, $h\in[0.5,0.9]$, $n_{\rm s}\in[0.8,1.2]$, $\sigma_{8}\in[0.6,1.0]$.
@@
We divide the simulation volume into $N^{3}$ cubic voxels and compute the density field in each voxel using a cloud in cell (CIC) estimator. For our fiducial analysis we used $N=128$, corresponding to voxels with a side length of $7.8h^{-1}{\rm Mpc}$, but we show that our results are insensitive to this choice (and thus to the smoothing scale of the density field) in Section 3.
@@ 3.1 LIMD halo biasing cannot reproduce the halo distribution
Figure 1: Power spectrum of the halo counts field before and after permuting voxels with approximately equal density, for halos of mass $10^{13.7}-10^{14.0}\,h^{-1}{\rm\,M}_{\sun}$ The permuted fields have significantly greater power on large scales, thus LIMD halo biasing cannot reproduce the distribution of halos. The mean and standard deviation are computed over 100 permuations. The vertical coloured lines correspond to the wavenumbers used to characterise this discrepancy in later plots.
@@
The vertical dashed lines in Fig. 1 correspond to four characteristic wavenumbers which we use for the remainder of the analysis to quantify the level of discrepancy, corresponding to approximately $0.01$, $0.03$, $0.1$ and $0.5\,h\,{\rm Mpc^{-1}}$. For our fiducial analysis, we find that $P_{\rm counts}(k)$ is a factor of $2.3\pm 0.3$, $1.6\pm 0.1$, $1.36\pm 0.03$ and $0.93\pm 0.01$ times the true $P_{\rm counts}(k)$, respectively, for these wavenumbers. It is therefore clear that the discrepancy is significant for this simulation, with the true power spectrum at approximately the $12\sigma$ value for $k=0.1\,h\,{\rm Mpc^{-1}}$.
@@
Figure 3: Fractional error on the power spectrum for halos of mass $10^{13.7}-10^{14.0}\,h^{-1}{\rm\,M}_{\sun}$ after permuting equal-density voxels as a function of voxel size. The black dashed line indicates the value of $N$ where we have as many voxels as density bins. For $N\geq 2^{5}$ we find significant disagreement on large scales, suggesting that one must smooth the density field over at least $60\,h^{-1}{\rm\,Mpc}$ for LIMD halo biasing to be valid. We do not plot values for wavenumbers beyond the Nyquist frequency, hence some lines are truncated.
Halo biasing is not just a function of halo mass, but also depends on the grid used to discretise the dark matter density field, or equivalently the smoothing scale of the field. In Fig. 3 we investigate how the error on $P_{\rm counts}(k)$ varies as a function of this grid size. One observes that the level of disagreement is relatively unchanged on the largest scales as one changes the voxel size from $\sim 4h^{-1}{\rm\,Mpc}$ to $\sim 30h^{-1}{\rm\,Mpc}$, indicating that LIMD biasing is inappropriate for density fields smoothed on these scales. As one further increases the voxel size to $\sim 60h^{-1}{\rm\,Mpc}$, the bias starts to diminish, with a bias of $\sim 4\sigma$ at $k=0.03\,h\,{\rm Mpc^{-1}}$. In this regime the number of voxels is similar to the number of density bins (as indicated by the black line in Fig. 3), but we have verified that similar behaviour is seen if we reduce $n_{\rm bin}$ to 100. Thus, the LIMD halo biasing is invalid for smoothed density fields with a smoothing scale smaller than $\sim 60h^{-1}{\rm\,Mpc}$, although it may be an acceptable approximation on larger scales.
@@ 3.2 Sensitivity to binning scheme
We have verified that our results are unchanged if we consider equally spaced bins in both density and the logarithm of density, instead of enforcing each bin to have approximately an equal number of voxels. The transient behaviour disappears at $n_{\rm bin}\approx 30$ when using logarithmic bins and at $n_{\rm bin}\approx 300$ for the linear binning scheme. Above these values, we again find that the bias is insensitive to the choice of $n_{\rm bin}$.
@@ 3.3 Sensitivity to cosmology and initial conditions
To test the robustness of our conclusion to the initial conditions, we repeated our analysis across the 100 high-resolution quijote simulations at the fiducial cosmology. We find biases (mean divided by standard deviation across 100 permutations) of $2.5\pm 1.0$, $7.9\pm 1.0$, $13.1\pm 1.2$ and $-8.7\pm 1.1$ at the wavenumbers shown in Fig. 1, indicating that our conclusion is insensitive to the initial white noise field.
Similarly, to investigate the sensitivity of the result to the cosmological parameters, we ran our analysis for all 2000 high-resolution simulations in the Latin hypercube suite. We find that, averaging over cosmological parameters, the biases for these wavenumbers are $2.9\pm 1.6$, $7.4\pm 1.4$, $12.8\pm 1.7$ and $-8.8\pm 1.7$, demonstrating that LIMD halo biasing does not hold for any of the cosmologies considered.
@@ 4 Discussion and conclusion
In this letter we have shown that the distribution of halos cannot be modelled according to LIMD biasing: the assumption that the number of halos in a given voxel is drawn from a probability distribution which depends solely on the local dark matter density and where each voxel is statistically independent. At least one of these assumptions must be broken to accurately model the distribution of halos in the Universe when averaging over scales of $\sim 4-60\,h^{-1}{\rm\,Mpc}$.
@@
This conclusion is particularly important as the use of higher-order statistics and field-level approaches become increasingly popular in cosmological analyses, since it is no longer sufficient to tune a bias model to only reproduce the power spectrum or other low-order statistics; one must be able to reproduce all moments of the field. In field-level studies, the initial conditions of the Universe are free parameters, so using a LIMD biasing model implies inferring an incorrect density field to obtain a predicted galaxy (or whichever tracer is used in the analysis) field which reasonably matches observations. Although the inferred galaxy field at the redshift of the observations would be trustworthy, extrapolating to other redshifts could be problematic and the the underlying density field and its initial conditions would have artefacts. Given that it is often these initial conditions which one is interested in, it is imperative that future studies move beyond the LIMD halo biasing assumption to avoid these problems.
@@
We conclude that number of halos in a given region of space is a stochastic function of the local dark matter density field, but it cannot depend solely on this quantity. Models which rely on this assumption, no matter how complex, cannot reproduce the distribution of halos and thus must be superseded by more accurate models. This letter not only disproves the LIMD assumption of halo biasing, but provides a framework to aid the design of future approaches by testing their assumptions in a model-independent manner.
Acknowledgements
We thank Simon Ding, Ludvig Doeser, Pedro Ferreira, Jens Jasche, Guilhem Lavaux, Stuart McAlpine and the Implicit Likelihood Inference, Accelerated Forward Modelling and Robustness working groups of the Simons Collaboration on “Learning the Universe” for useful comments and suggestions.
DJB and MH are supported by the Simons Collaboration on “Learning the Universe.” BDW acknowleges the DIM-ORIGINES-2023, Infinity Next grant. The Flatiron Institute is supported by the Simons Foundation.
For the purposes of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.

### v2

# Bye bye, local bias: the statistics of the halo field are poorly determined by the local mass density

@@ Abstract
Bias models relating the dark matter field to the spatial distribution of halos are widely used in current cosmological analyses. Many models predict halos purely from the local Eulerian matter density, yet bias models in perturbation theory require the inclusion of other local properties. We assess the validity of assuming that only the local dark matter density can be used to predict the number density of halos in a model-independent way and in the non-perturbative regime. Utilising $N$-body simulations, we study the properties of the halo counts field after spatial voxels with near-equal dark matter density have been permuted. If local-in-matter-density biasing were valid, the statistical properties of the permuted and un-permuted fields would be indistinguishable since both represent equally fair draws of the stochastic biasing model. For voxels of side length $\sim 4-30\,h^{-1}{\rm\,Mpc}$ and for halos less massive than $\sim 10^{15}\,h^{-1}{\rm\,M_{\sun}}$, we find that the permuted halo field has a scale-dependent bias with greater than 25% more power on scales relevant for current surveys. These bias models remove small-scale power by not modelling correlations between neighbouring voxels, which substantially boosts large-scale power to conserve the field’s total variance. This conclusion is robust to the choice of initial conditions and cosmology. Assuming local-in-matter-density halo biasing cannot, therefore, reproduce the distribution of halos across a large range of scales and halo masses, no matter how complex the model. One must either allow the biasing to be a function of other quantities and/or remove the assumption that neighbouring voxels are statistically independent.
Cosmology (343); Dark marrer distribution (356); Large-scale structure of the universe (902)
journal: ApJLsoftware: astropy (Astropy Collaboration et al., 2013, 2018, 2022), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), pandas (Wes McKinney, 2010; pandas development team, 2020), pylians (Villaescusa-Navarro, 2018).
@@ 1 Introduction
On the largest scales, one expects that the bias is linear, namely the overdensity of tracers is proportional to the overdensity of matter. This approximation breaks down as one moves from linear to quasilinear scales, and more general functional relationships must be used. The simplest assumption one can make is that the tracer number density depends solely on the local Eulerian dark matter density - the so-called “local-in-matter-density” (LIMD) assumption (Desjacques et al., 2018) which, in a perturbative analysis, allows one to write the tracer number density as an expansion in the overdensity: the local bias expansion (Fry & Gaztanaga, 1993). More generally, one can express the bias model as local in gravitational observables, namely not just the density field but other terms related to the tidal field or higher derivative terms (Mirbabayi et al., 2015; Desjacques et al., 2018; Lazeyras & Schmidt, 2018, 2019). One can enumerate such terms based on the order of perturbation theory such that, if one goes to sufficiently high order and is in the large-scale, perturbative regime, then the result converges to the truth.
@@
Relying on perturbation theory (Bernardeau et al., 2002) or effective field theory (EFT) (Carrasco et al., 2012; Senatore & Zaldarriaga, 2014; Senatore, 2015; Perko et al., 2016) inherently limits our analysis to these quasi-linear scales. Although it is possible to study higher order statistics (D’Amico et al., 2022) or perform field-level analyses in these frameworks (Schmidt et al., 2019; Elsner et al., 2020; Kostić et al., 2023; Stadler et al., 2023; Nguyen et al., 2024), to extract information on non-linear scales one may need to go beyond such local bias expansions, although some studies have shown that perturbative methods perform well when tested against $N$-body simulations (Roth & Porciani, 2011; Schmittfull et al., 2019). Instead of relying of power series, many parametric forms for the biasing relation have been proposed in the literature based on the LIMD assumption (Szalay, 1988; Matsubara, 1995, 2011; Frusciante & Sheth, 2012; Neyrinck, 2014; Ata et al., 2015). As such, many reconstruction analyses typically assume that the number density of halos is solely a function of the smoothed density field (e.g. Schmoldt et al., 1999; Erdoǧdu et al., 2004, 2006; Kitaura et al., 2009; Jasche et al., 2010; Jasche & Wandelt, 2012, 2013; Jasche et al., 2015; Lavaux & Jasche, 2016; Modi et al., 2018; Jasche & Lavaux, 2019; Lavaux et al., 2019; Ramanah et al., 2019). Alternatively, one can utilise machine learning methods to map the dark matter density field (using both local and non-local information) to the halo counts field (Charnock et al. (2020); see Dai & Seljak (2021) for a Lagrangian approach).
@@
Unlike in the perturbative regime, we do not have theoretical guarantees that any of these bias models are accurate and so this must be tested empirically. Mirbabayi et al. (2015) argue that, at any order in perturbation theory, the bias can be expressed in terms of locally measurable observables, and that this argument is closed under renormalisation. However, it could be (as is often assumed (Szalay, 1988; Matsubara, 1995, 2011; Frusciante & Sheth, 2012; Neyrinck, 2014; Ata et al., 2015)) that the only important observable is the local dark matter density field. Testing this assumption is the subject of this letter. There is an important distinction, however, between testing the assumptions of the halo biasing model (i.e. which variables are needed to predict the tracer number density) and the particular model chosen (i.e. what is the functional form of the bias or the stochasticity in the tracer-matter relation). Previous tests of the latter seem to suggest that information beyond the local density is required (e.g. Nguyen et al., 2021; Fang et al., 2024), however this could just be due to insufficiently accurate LIMD biasing models. Initial studies of the former also seem to support this conclusion (Balaguera-Antolínez et al., 2019; Pellejero-Ibañez et al., 2020; Balaguera-Antolínez et al., 2020).
In this letter we study the former issue in the fully non-linear regime.
@@
We present a model-independent test of the assumption of LIMD halo biasing. Our test is independent of the functional form of the biasing model, but simply tests whether halo biasing can be a stochastic function of only the local matter density. We find that LIMD halo biasing cannot reproduce the distribution of halos less massive than $\sim 10^{15}\,h^{-1}\,{\rm M_{\sun}}$ when applied to gridded dark matter fields with voxels of size less than $\sim 30\,h^{-1}{\rm\,Mpc}$. Our permutation test naturally enforces conservation of all 1-point moments and count statistics, while demonstrating the insufficiency of the LIMD hyopthesis to model halo clustering. We demonstrate that this conclusion is robust across initial conditions and cosmology by utilising the quijote suite of $N$-body simulations. We conclude that no LIMD biasing model, however complex, should be used if one wishes to accurately predict the spatial distribution of halos given a matter density field.
@@
In Sec. 2 we explicitly define LIMD biasing, introduce our model-independent test and describe the simulations used to perform this test. Our results are presented in Sec. 3, where we demonstrate that our conclusions are highly robust to the one free parameter in our test, as well as cosmology and initial conditions, and investigate the adequacy of LIMD biasing as a function of scale and halo mass. Sec. 4 discusses these results and presents our conclusions.
@@ 2.1 Assessing the assumption of local biasing
Since the dark matter density field takes continuous values, instead of performing permutations on voxels which have exactly equal dark matter density, we first sort the voxels into $n_{\rm bin}$ bins of dark matter density, and consider two voxels equivalent if they fall within the same bin, namely if their density is approximately equal. In our fiducial analysis, we choose to use adaptive bin widths so that each bin has approximately the same number of voxels and use $n_{\rm bin}=1000$. We demonstrate that our results are highly robust to this choice in Sec. 3.
@@ 2.2 Simulations
For our fiducial analysis, we consider simulation number 0 from the fiducial cosmology, which matches the Planck 2018 cosmological parameters (Planck Collaboration, 2020): $\Omega_{\rm m}=0.3175$, $\Omega_{\rm b}=0.049$, $h=0.6711$, $n_{\rm s}=0.9624$, $\sigma_{8}=0.834$. In Sec. 3 we verify that our conclusions are robust if we chose different initial conditions (i.e. for different simulation numbers in the suite). We also study the sensitivity of our conclusions to the choice of cosmological parameters by performing our analysis with the (high-resolution) Latin-hypercube suite of quijote simulations, consisting of 2000 dark-matter-only simulations with cosmological parameters arranged on a Latin hypercube in the range: $\Omega_{\rm m}\in[0.1,0.5]$, $\Omega_{\rm b}\in[0.03,0.07]$, $h\in[0.5,0.9]$, $n_{\rm s}\in[0.8,1.2]$, $\sigma_{8}\in[0.6,1.0]$.
@@
We divide the simulation volume into $N^{3}$ cubic voxels and compute the density field in each voxel using a nearest grid point (NGP) estimator. This is the same estimator used to compute the gridded halo counts field, and thus we mitigate effects which arise due to different discretisation procedures. We have tested our method using a range of other density field estimators: cloud in cell (CIC), triangular-shape cloud (TSC), and piecewise cubic spline (PCS) kernel, as well as one inspired by smoothed particle hydrodynamics (SPH) (Monaghan, 1992; Colombi et al., 2007). For our fiducial analysis we used $N=128$, corresponding to voxels with a side length of $\Delta x=7.8h^{-1}{\rm Mpc}$, but we show that our results are insensitive to this choice (and thus to the smoothing scale of the density field) in Sec. 3. At our fiducial grid resolution, we find that all density estimators give the same qualitative trends, although the magnitude of the bias induced by LIMD models varies with this choice. For large voxel sizes, we find the performance is best for the NGP density estimator, but the degree of bias depends sensitively on the choice of mass assignment in the definition of the density field and is much worse for other choices. We thus choose the NGP estimator to ensure our results are conservative.
@@ 3.1 LIMD halo biasing cannot reproduce the halo distribution
Figure 1: Power spectrum of the halo counts field before and after permuting voxels with approximately equal density, for halos of mass $10^{13.7}-10^{14.0}\,h^{-1}{\rm\,M}_{\sun}$ The permuted fields have significantly greater power on large scales, thus LIMD halo biasing cannot reproduce the distribution of halos. The mean and standard deviation are computed over 100 permutations. The vertical coloured lines correspond to the wavenumbers used to characterise this discrepancy in later plots.
@@
The vertical dashed lines in Fig. 1 correspond to four characteristic wavenumbers which we use for the remainder of the analysis to quantify the level of discrepancy, corresponding to approximately $0.01$, $0.03$, $0.1$ and $0.5\,h\,{\rm Mpc^{-1}}$. For our fiducial analysis, we find that $P_{\rm counts}(k)$ is a factor of $1.6\pm 0.2$, $1.3\pm 0.07$, $1.16\pm 0.02$ and $0.97\pm 0.01$ times the true $P_{\rm counts}(k)$, respectively, for these wavenumbers. It is therefore clear that the discrepancy is significant for this simulation, with the true power spectrum at approximately the $7.6\sigma$ value for $k=0.1\,h\,{\rm Mpc^{-1}}$.
@@
Figure 3: Fractional error on the power spectrum for halos of mass $10^{13.7}-10^{14.0}\,h^{-1}{\rm\,M}_{\sun}$ after permuting equal-density voxels as a function of voxel size. The black dashed line indicates the value of $N$ where we have as many voxels as density bins. For $N\geq 2^{6}$ we find significant disagreement on large scales, suggesting that one must smooth the density field over at least $30\,h^{-1}{\rm\,Mpc}$ for LIMD halo biasing to be valid. We do not plot values for wavenumbers beyond the Nyquist frequency, hence some lines are truncated.
[…]
@@ 3.2 Sensitivity to binning scheme
We have verified that our results are unchanged if we consider equally spaced bins in both density and the logarithm of density, instead of enforcing each bin to have approximately an equal number of voxels. The transient behaviour disappears at $n_{\rm bin}\approx 10$ when using logarithmic bins and at $n_{\rm bin}\approx 300$ for the linear binning scheme. Above these values, we again find that the bias is insensitive to the choice of $n_{\rm bin}$.
@@ 3.3 Sensitivity to cosmology and initial conditions
To test the robustness of our conclusion to the initial conditions, we repeated our analysis across the 100 high-resolution quijote simulations at the fiducial cosmology. We find biases (mean divided by standard deviation across 100 permutations) of $1.4\pm 0.9$, $4.5\pm 1.0$, $6.6\pm 0.9$, $-4.7\pm 1.0$ at the wavenumbers shown in Fig. 1, indicating that our conclusion is insensitive to the initial white noise field.
Similarly, to investigate the sensitivity of the result to the cosmological parameters, we ran our analysis for all 2000 high-resolution simulations in the Latin hypercube suite. We find that, averaging over cosmological parameters, the biases for these wavenumbers are $1.7\pm 1.4$, $4.2\pm 1.5$, $6.4\pm 1.2$, $-4.5\pm 1.1$ demonstrating that LIMD halo biasing does not hold for any of the cosmologies considered.
@@ 4 Discussion and conclusion
In this letter we have shown that the distribution of halos cannot be modelled according to LIMD biasing: the assumption that the number of halos in a given voxel is drawn from a probability distribution which depends solely on the local dark matter density and where each voxel is statistically independent. At least one of these assumptions must be broken to accurately model the distribution of halos in the Universe when averaging over scales of $\sim 4-30\,h^{-1}{\rm\,Mpc}$.
@@
The test described in this letter is similar in approach to the first step of the Bias Assignment Method (BAM) (Balaguera-Antolínez et al., 2019), which is used to fit non-parametric bias models. As here, Balaguera-Antolínez et al. (2019) find that information beyond the local density is required to recover the halo distribution but perform their analysis with a single grid resolution of $\Delta x=3\,h^{-1}{\rm\,Mpc}$, which is slightly finer than we consider. Our work extends this conclusion by systematically exploring the sensitivity to this smoothing scale and explicitly investigating the mass dependence of this conclusion in separate mass bins. BAM has recently been used to provide evidence for non-locality in the bias model by studying the bispectrum (Pellejero-Ibañez et al., 2020), with up to $4.8\sigma$ significance (Kitaura et al., 2022). The conclusions presented here are much stronger than this. Moreover, we demonstrate that the results are robust to cosmological parameters and provide more thorough tests of the robustness of our results to the binning scheme. Unlike Balaguera-Antolínez et al. (2019), we have demonstrated that our conclusions are robust to the resolution of the $N$-body simulation used in our test.
This conclusion is particularly important as the use of higher-order statistics and field-level approaches become increasingly popular in cosmological analyses, since it is no longer sufficient to tune a bias model to only reproduce the power spectrum or other low-order statistics; one must be able to reproduce all moments of the field. In field-level studies, the initial conditions of the Universe are free parameters, so using a LIMD biasing model implies inferring an incorrect density field to obtain a predicted galaxy (or whichever tracer is used in the analysis) field which reasonably matches observations (Nguyen et al., 2021). Although the inferred galaxy field at the redshift of the observations would be trustworthy, extrapolating to other redshifts could be problematic and the underlying density field and its initial conditions would have artefacts. Given that it is often these initial conditions which one is interested in, it is imperative that future studies move beyond the LIMD halo biasing assumption to avoid these problems.
@@
As discussed in Sec. 2, the degree to which LIMD biasing fails depends on the estimator used to obtain the density field. This high sensitivity to the estimator is problematic if one wishes to predict the halo distribution from an approximate gravity model. As such, using information beyond the local density field is essential to produce realistic halo catalogues for analysis pipelines for current and future surveys.
We conclude that number of halos in a given region of space is a stochastic function of the local dark matter density field, but it cannot depend solely on this quantity. Models which rely on this assumption, no matter how complex, cannot reproduce the distribution of halos and thus must be superseded by more accurate models. This letter not only disproves the LIMD assumption of halo biasing for scales $\lesssim 30\,h^{-1}{\rm\,Mpc}$, but provides a framework to aid the design of future approaches by testing their assumptions in a model-independent manner.
We thank Andrés Balaguera, Yan-Chuan Cai, Simon Ding, Ludvig Doeser, Rozo Eduardo, Pedro Ferreira, Joachim Harnois-Deraps, Jens Jasche, Guilhem Lavaux, Titouan Lazeyras, Stuart McAlpine, Minh Nguyen, Francisco Villaescusa-Navarro and the Implicit Likelihood Inference, Accelerated Forward Modelling and Robustness working groups of the Simons Collaboration on “Learning the Universe” for useful comments and suggestions. We particularly thank Neal Dalal and Cora Uhlemann for independently checking our analysis after it appeared as a pre-print, and prompting us to consider the sensitivity of our results to the mass assignment scheme. DJB and MH are supported by the Simons Collaboration on “Learning the Universe.” BDW acknowleges the DIM-ORIGINES-2023, Infinity Next grant. The Flatiron Institute is supported by the Simons Foundation. This work was done within the Aquila Consortium (https://www.aquila-consortium.org/). For the purposes of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.